# Lossy Beyond Diagonal Reconfigurable Intelligent Surfaces: Modeling and Optimization

### v1

@@ III-B.1 MM Approximation
where we extract all entries of interest in $\mathbf{\Phi}$ to construct $\bar{\bm{\phi}}=[\bar{\bm{\phi}}_{1}^{T},\ldots,\bar{\bm{\phi}}_{G}^{T}]^{T}$ with each $\bar{\bm{\phi}}_{g}\in\mathbb{C}^{\frac{\bar{M}(\bar{M}+1)}{2}\times 1}$ containing the diagonal and upper-triangular entries of $\mathbf{\Phi}_{g}$, i.e., $\bm{\phi}_{g}=\mathsf{vec}(\mathbf{\Phi}_{g})=\mathbf{P}\bar{\bm{\phi}}_{g}$, and $\bar{\mathbf{f}}_{t}$ is a given vector constructed based on $\mathbf{F}$, $\tilde{\bm{\phi}}_{t}$, $\boldsymbol{l}$, and $\mathbf{P}$ to align with the structure of $\bar{\bm{\phi}}$. Defining $\tilde{\mathbf{f}}_{t}\triangleq\mathbf{F}\tilde{\bm{\phi}}_{t}+\boldsymbol{l}$, $\tilde{\mathbf{F}}_{t}\triangleq\overline{\mathsf{vec}}(\tilde{\mathbf{f}}_{t})$, and $\mathbf{P}_{G}\triangleq\mathbf{I}_{G}\otimes\mathbf{P}$, we can prove the transformation in (20a) as follows:
@@
$=\begin{bmatrix}\mathsf{vec}^{H}(\tilde{\mathbf{F}}_{t,1}),&\ldots,&\mathsf{vec}^{H}(\tilde{\mathbf{F}}_{t,G})\end{bmatrix}\mathbf{P}_{G}\bar{\bm{\phi}},$ (21)
where each $\tilde{\mathbf{F}}_{t,g}$ represents a submatrix of $\tilde{\mathbf{F}}_{t}$, specifically denoted as $[\tilde{\mathbf{F}}_{t}]_{(g-1)\bar{M}+1:g\bar{M},(g-1)\bar{M}+1:g\bar{M}}$. Defining $\bar{\mathbf{f}}_{t}$ as $\bar{\mathbf{f}}_{t}^{H}\triangleq\begin{bmatrix}\mathsf{vec}^{H}(\tilde{\mathbf{F}}_{t,1}),&\ldots,&\mathsf{vec}^{H}(\tilde{\mathbf{F}}_{t,G})\end{bmatrix}\mathbf{P}_{G}$, then we have the equality $(\mathbf{F}\tilde{\bm{\phi}}_{t}+\boldsymbol{l})^{H}\tilde{\bm{\phi}}=\bar{\mathbf{f}}_{t}^{H}\bar{\bm{\phi}}$.
@@ III-B.2 ADMM Methodology
Following the MM approximation in the outer loop, the ADMM method is applied in the inner loop to deal with the nonlinear coupling constraint (18b). We first note that when the variable $\mathbf{Y}_{g}/\bm{\Phi}_{g}$ is fixed, constraint (18b) is linear in $\bm{\Phi}_{g}/\mathbf{Y}_{g}$, as written below:
$(Y_{0}\mathbf{I}_{\bar{M}}+\mathbf{Y}_{g})\bm{\Phi}_{g}$ $=Y_{0}\mathbf{I}_{\bar{M}}-\mathbf{Y}_{g},\forall g\in\mathcal{G},$ (22)
$\mathbf{Y}_{g}(\bm{\Phi}_{g}+\mathbf{I}_{\bar{M}})$ $=Y_{0}(\mathbf{I}_{\bar{M}}-\bm{\Phi}_{g}),\forall g\in\mathcal{G}.$ (23)
@@
$\mathbf{C}_{g}(\bar{\mathbf{y}}_{g})\bar{\bm{\phi}}_{g}$ $=\mathbf{c}_{g}(\bar{\mathbf{y}}_{g}),\forall g\in\mathcal{G},$ (24)
$\mathbf{D}_{g}(\bar{\bm{\phi}}_{g})\bar{\mathbf{y}}_{g}$ $=\mathbf{d}_{g}(\bar{\bm{\phi}}_{g}),\forall g\in\mathcal{G},$ (25)
where $\mathbf{C}_{g}(\bar{\mathbf{y}}_{g})\triangleq(\mathbf{I}_{\bar{M}}\otimes(Y_{0}\mathbf{I}_{\bar{M}}+\overline{\mathsf{vec}}(\mathbf{B}\mathbf{P}\bar{\mathbf{y}}_{g})))\mathbf{P}$, $\mathbf{c}_{g}(\bar{\mathbf{y}}_{g})\triangleq\mathsf{vec}(Y_{0}\mathbf{I}_{\bar{M}}-\overline{\mathsf{vec}}(\mathbf{B}\mathbf{P}\bar{\mathbf{y}}_{g}))$, $\mathbf{D}_{g}(\bar{\bm{\phi}}_{g})\triangleq(((\overline{\mathsf{vec}}(\mathbf{P}\bar{\bm{\phi}}_{g})+\mathbf{I}_{\bar{M}})^{T}\otimes\mathbf{I}_{\bar{M}})\mathbf{B}\mathbf{P}$, and $\mathbf{d}_{g}(\bar{\bm{\phi}}_{g})\triangleq\mathsf{vec}(Y_{0}\mathbf{I}_{\bar{M}}-\overline{\mathsf{vec}}(\mathbf{P}\bar{\bm{\phi}}_{g}))$. In the following ADMM framework, (24) will be employed for updating the variable $\bar{\bm{\phi}}$, while (25) will be used for updating the variable $\bar{\mathbf{y}}_{g}$. However, updating $\bar{\mathbf{y}}_{g}$ meets a challenge due to the non-separable nature between entries of $\bar{\mathbf{y}}_{g}$ in (25), which makes it impossible to individually optimize each entry in $\bar{\mathbf{y}}_{g}$ under the non-convex constraint (18d). To overcome this, one effective solution is to introduce an auxiliary variable $\mathbf{z}$ by setting $\mathbf{z}=\bar{\mathbf{y}}$, where $\bar{\mathbf{y}}=[\bar{\mathbf{y}}_{1}^{T},\ldots,\bar{\mathbf{y}}_{G}^{T}]^{T}$, thereby decoupling (25) from the troublesome constraints (18d). Hence, problem (20) is equivalently re-written as
@@
$[\mathbf{z}]_{\tilde{m}}\in\mathcal{Y},\forall\tilde{m}\in\tilde{\mathcal{M}},$ (26d)
where ~=$\tilde{\mathcal{M}}=\{1,\ldots,\frac{M(\bar{M}+1)}{2}\}$. Following the ADMM framework, we construct the augmented Lagrangian (AL) function in its scaled form [29], given by
@@
where $\mathbf{C}=\mathsf{blkdiag}(\mathbf{C}_{1},\ldots,\mathbf{C}_{G})$, $\mathbf{c}=[\mathbf{c}_{1}^{T},\ldots,\mathbf{c}_{G}^{T}]^{T}$, and $\bm{\lambda}_{1}\in\mathbb{C}^{G\bar{M}^{2}\times 1}$, $\bm{\lambda}_{2}\in\mathbb{C}^{\frac{M(\bar{M}+1)}{2}\times 1}$ denote the scaled Lagrange multipliers corresponding to constraints (26b) and (26c), respectively, and $\rho_{1}$ and $\rho_{2}$ are penalty coefficients.
@@
    $\bar{\bm{\phi}}=\left(\rho_{1}\mathbf{C}^{H}\mathbf{C}\right)^{-1}\left(2\bar{\mathbf{f}}_{t}+\rho_{1}\mathbf{C}^{H}(\mathbf{c}-\bm{\lambda}_{1})\right);$ (29)
@@
    $\bar{\mathbf{y}}=\Big{(}\rho_{1}\mathbf{D}^{H}\mathbf{D}$ $+\rho_{2}\mathbf{I}_{M(\bar{M}+1)/2}\Big{)}^{-1}$ (31)
    ×$\times\left(\rho_{1}\mathbf{D}^{H}(\mathbf{d}-\bm{\lambda}_{1})+\rho_{2}(\mathbf{z}+\bm{\lambda}_{2})\right);$
@@
    Since constraint (26d) is separable with respect to each entry of vector $\mathbf{z}$, problem (32) can be decomposed into $\frac{M(\bar{M}+1)}{2}$ independent one-dimensional problems. Defining $\mathbf{r}\triangleq\bar{\mathbf{y}}-\bm{\lambda}_{2}$, each one-dimensional problem can be expressed as
@@
    $\theta_{\tilde{m}}=\begin{cases}\theta_{\mathrm{min}},&\text{if}~{}\arg\psi\leq\theta_{\mathrm{min}};\\ \theta_{\mathrm{max}},&\text{if}~{}\arg\psi\geq\theta_{\mathrm{max}};\\ \arg\psi,&\text{if}~{}\theta_{\mathrm{min}}<\arg\psi<\theta_{\mathrm{max}};\end{cases}$ (35)
@@
Algorithm 1 Proposed MM-ADMM Algorithm
0:  $h_{RT}$, $\mathbf{h}_{RI}$, $\mathbf{h}_{IT}$, $\alpha$, $\beta$, $\bar{M}$, $\theta_{\mathrm{min}}$, $\theta_{\mathrm{max}}$.
0:  $\bar{\bm{\phi}}^{\star}$.
@@
4:  while no convergence of objective (18ado
5:     while no convergence of objective (20ado
@@
8:        Update $[\mathbf{z}]_{\tilde{m}},\forall\tilde{m}\in\tilde{\mathcal{M}}$ by (35).
@@ III-B.3 Summary
The proposed MM-ADMM algorithm is summarized in Algorithm 1. Once $\bar{\bm{\phi}}^{\star}$ is obtained, we can compute $\bm{\Phi}^{\star}_{g}$ as $\bm{\Phi}^{\star}_{g}=\overline{\mathsf{vec}}(\mathbf{P}\bar{\bm{\phi}}^{\star}_{g}),\forall g\in\mathcal{G}$. The final solution for $\bm{\Phi}^{\star}$ is then given by $\bm{\Phi}^{\star}=\mathsf{blkdiag}(\bm{\Phi}^{\star}_{1},\ldots,\bm{\Phi}^{\star}_{G})$.
@@ III-C Low-Complexity Algorithm
The proposed MM-ADMM algorithm requires a double-loop iteration, which might require high computational complexity to guarantee convergence. To ease the requirement of the double-loop iteration, we first recall that, for a lossless group-connected reconfigurable admittance network, the received signal power $\left|h_{RT}+\mathbf{h}_{RI}^{H}\bm{\Phi}\mathbf{h}_{IT}\right|^{2}$ reaches its upper bound when $\bm{\Phi}$ satisfies [5]
$e^{\jmath\arg(h_{RT})}\bar{\mathbf{h}}_{RI,g}=\bm{\Phi}_{g}\bar{\mathbf{h}}_{IT,g},\forall g\in\mathcal{G},$ (38)
where $\bar{\mathbf{h}}_{RI}=[\bar{\mathbf{h}}_{RI,1}^{T},\ldots,\bar{\mathbf{h}}_{RI,G}^{T}]^{T}$ with $\bar{\mathbf{h}}_{RI,g}\triangleq\frac{\mathbf{h}_{RI,g}}{\|\mathbf{h}_{RI,g}\|_{2}}\in\mathbb{C}^{\bar{M}\times 1}$, and $\bar{\mathbf{h}}_{IT}=[\bar{\mathbf{h}}_{IT,1}^{T},\ldots,\bar{\mathbf{h}}_{IT,G}^{T}]^{T}$ with $\bar{\mathbf{h}}_{IT,g}\triangleq\frac{\mathbf{h}_{IT,g}}{\|\mathbf{h}_{IT,g}\|_{2}}\in\mathbb{C}^{\bar{M}\times 1}$. By substituting $\bm{\Phi}_{g}$ with $\mathbf{Y}_{g},\forall g\in\mathcal{G}$ based on (18b), (38) can be equivalently re-written as
@@
$\mathrm{s.t.}~{}~{}$ $[\bar{\mathbf{y}}]_{\tilde{m}}\in\mathcal{Y},\forall\tilde{m}\in\tilde{\mathcal{M}},$ (41b)
[…]
@@
$\mathbf{Y}_{g}\mathbf{a}_{g}=\mathbf{A}_{g}\bar{\mathbf{y}}_{g},\forall g\in\mathcal{G},$ (43)
@@
We apply the ADMM method [29] to solve (44). The AL function of (44) is given by:
@@
where $\mathbf{u}\in\mathbb{C}^{\frac{M(\bar{M}+1)}{2}\times 1}$ is the scaled Lagrangian multiplier and $\rho$ is the penalty coefficient. The detailed algorithm is presented below.
@@
    $\bar{\mathbf{y}}=\left(2\mathbf{A}^{H}\mathbf{A}+\rho\mathbf{I}_{M(\bar{M}+1)/2}\right)^{-1}\left(2\mathbf{A}^{H}\mathbf{b}+\rho(\mathbf{z}+\mathbf{u})\right);$ (47)
@@
With the solution $\bar{\mathbf{y}}^{\star}$ to problem (44), we can compute $\mathbf{Y}_{g}^{\star},\forall g\in\mathcal{G}$ as $\mathbf{Y}_{g}^{\star}=\overline{\mathsf{vec}}(\mathbf{B}\mathbf{P}\bar{\mathbf{y}}_{g}^{\star})$. The admittance matrix is then given by $\mathbf{Y}^{\star}=\mathsf{blkdiag}(\mathbf{Y}_{1}^{\star},\ldots,\mathbf{Y}_{G}^{\star})$, from which the scattering matrix $\bm{\Phi}^{\star}$ is obtained using (1).
III-D Extension to Forest-Connected BD-RIS
The proposed optimization algorithms above are formulated for the group-connected BD-RIS. To adapt these algorithms for the forest-connected BD-RIS, the modification is straightforward: the mapping $\bar{\mathbf{Y}}_{g}=\overline{\mathsf{vec}}(\mathbf{P}\bar{\mathbf{y}}_{g})$, originally designed for the group-connected architecture, needs to be adjusted to align with the forest-connected architecture. For the tridiagonal form of the tree-connected architecture of each $\bar{\mathbf{Y}}_{g}$, (12d), (12e) and (5) show that $\bar{\mathbf{Y}}_{g},\forall g\in\mathcal{G}$ is determined by its $\bar{M}$ diagonal entries and $\bar{M}-1$ off-diagonal entries specified by (5). Similarly, for the arrowhead form, each $\bar{\mathbf{Y}}_{g}$ is determined by its $\bar{M}$ diagonal entries and $\bar{M}-1$ off-diagonal entries specified by (6). This inspires the extraction of $12\bar{M}-$ 1 entries from each tridiagonal $\bar{\mathbf{Y}}_{g}$ to construct $\tilde{{\mathbf{y}}}_{g}\in\mathbb{C}^{(2\bar{M}-1)\times 1}$. To achieve this, we define $\bar{\mathbf{Y}}_{g}=\overline{\mathsf{vec}}(\tilde{\mathbf{P}}\tilde{{\mathbf{y}}}_{g})$. The mapping matrix $\tilde{\mathbf{P}}\in\{0,1\}^{\bar{M}^{2}\times(2\bar{M}-1)}$ is specified as follows:
$[\tilde{\mathbf{P}}]_{\bar{M}(i-1)+j,k}=$ (50)
$\begin{cases}1,&k=j,~{}i=1,~{}\text{and}~{}j\leq 2;\\ 1,&k=\bar{M}+j-1,~{}i=\bar{M}~{}\text{and}~{}j\geq i-1;\\ 1,&k=i-1+j,~{}1<i<\bar{M}~{}\text{and}~{}i-1\leq j\leq i+1;\\ 0,&\text{otherwise},\end{cases}$
[…]
$[\breve{\mathbf{P}}]_{\bar{M}(i-1)+j,k}=\begin{cases}1,&k=j,~{}\text{and}~{}i=1;\\ 1,&k=i,~{}j=1~{}\text{and}~{}i>1;\\ 1,&k=\bar{M}+i-1,~{}i=i~{}\text{and}~{}i>1;\\ 0,&\text{otherwise},\end{cases}$ (51)
[…]
III-E Complexity Analysis

### v2

@@ III-B.1 $M$ Approximation
where we extract all entries of interest in $\mathbf{\Phi}$ to construct $\bar{\bm{\phi}}=[\bar{\bm{\phi}}_{1}^{T},\ldots,\bar{\bm{\phi}}_{G}^{T}]^{T}$ with each $\bar{\bm{\phi}}_{g}\in\mathbb{C}^{\frac{\bar{M}(\bar{M}+1)}{2}\times 1}$ containing the diagonal and upper-triangular entries of $\mathbf{\Phi}_{g}$, i.e., $\bm{\phi}_{g}=\mathsf{vec}(\mathbf{\Phi}_{g})=\mathbf{P}\bar{\bm{\phi}}_{g},~\forall g\in\mathcal{G}$. The matrix $\mathbf{P}$ used to extract the diagonal and upper-triangular entries of $\mathbf{\Phi}_{g}$ is the same as defined in (62). $\bar{\mathbf{f}}_{t}$ is a given vector constructed based on $\mathbf{F}$, $\tilde{\bm{\phi}}_{t}$, $\boldsymbol{l}$, and $\mathbf{P}$ to align with the structure of $\bar{\bm{\phi}}$. Defining $\tilde{\mathbf{f}}_{t}\triangleq\mathbf{F}\tilde{\bm{\phi}}_{t}+\boldsymbol{l}$, $\tilde{\mathbf{F}}_{t}\triangleq\overline{\mathsf{vec}}(\tilde{\mathbf{f}}_{t})$, and $\mathbf{P}_{G}\triangleq\mathbf{I}_{G}\otimes\mathbf{P}$, we can prove the transformation in (19a) as follows:
@@
$=\begin{bmatrix}\mathsf{vec}^{H}(\tilde{\mathbf{F}}_{t,1}),\ldots,\mathsf{vec}^{H}(\tilde{\mathbf{F}}_{t,G})\end{bmatrix}\mathbf{P}_{G}\bar{\bm{\phi}},$ (20)
where each $\tilde{\mathbf{F}}_{t,g}$ represents a submatrix of $\tilde{\mathbf{F}}_{t}$, specifically denoted as $\tilde{\mathbf{F}}_{t}]_{(g-1)\bar{M}+1:g\bar{M},(g-1)\bar{M}+1:g\bar{M}}$. Defining $\bar{\mathbf{f}}_{t}$ as $\bar{\mathbf{f}}_{t}^{H}\triangleq\begin{bmatrix}\mathsf{vec}^{H}(\tilde{\mathbf{F}}_{t,1}),\ldots,\mathsf{vec}^{H}(\tilde{\mathbf{F}}_{t,G})\end{bmatrix}\mathbf{P}_{G}$, then we have the equality $(\mathbf{F}\tilde{\bm{\phi}}_{t}+\boldsymbol{l})^{H}\tilde{\bm{\phi}}=\bar{\mathbf{f}}_{t}^{H}\bar{\bm{\phi}}$.
@@ III-B.2 ADMM Methodology
Following the $M$ approximation in the outer loop, the ADMM method is applied in the inner loop to deal with the nonlinear coupling constraint (17b). We first note that when the variable $\mathbf{Y}_{g}/\bm{\Phi}_{g}$ is fixed, constraint (17b) is linear in $\bm{\Phi}_{g}/\mathbf{Y}_{g}$, as written below:
$(Y_{0}\mathbf{I}_{\bar{M}}+\mathbf{Y}_{g})\bm{\Phi}_{g}$ $=Y_{0}\mathbf{I}_{\bar{M}}-\mathbf{Y}_{g},~\forall g\in\mathcal{G},$ (21)
$\mathbf{Y}_{g}(\bm{\Phi}_{g}+\mathbf{I}_{\bar{M}})$ $=Y_{0}(\mathbf{I}_{\bar{M}}-\bm{\Phi}_{g}),~\forall g\in\mathcal{G}.$ (22)
@@
$\mathbf{C}_{g}(\bar{\mathbf{y}}_{g})\bar{\bm{\phi}}_{g}$ $=\mathbf{c}_{g}(\bar{\mathbf{y}}_{g}),~\forall g\in\mathcal{G},$ (23)
$\mathbf{D}_{g}(\bar{\bm{\phi}}_{g})\bar{\mathbf{y}}_{g}$ $=\mathbf{d}_{g}(\bar{\bm{\phi}}_{g}),~\forall g\in\mathcal{G},$ (24)
where $\mathbf{C}_{g}(\bar{\mathbf{y}}_{g})\triangleq(\mathbf{I}_{\bar{M}}\otimes(Y_{0}\mathbf{I}_{\bar{M}}+\overline{\mathsf{vec}}(\mathbf{B}\mathbf{P}\bar{\mathbf{y}}_{g})))\mathbf{P}$, $\mathbf{c}_{g}(\bar{\mathbf{y}}_{g})\triangleq\mathsf{vec}(Y_{0}\mathbf{I}_{\bar{M}}-\overline{\mathsf{vec}}(\mathbf{B}\mathbf{P}\bar{\mathbf{y}}_{g}))$, $\mathbf{D}_{g}(\bar{\bm{\phi}}_{g})\triangleq(((\overline{\mathsf{vec}}(\mathbf{P}\bar{\bm{\phi}}_{g})+\mathbf{I}_{\bar{M}})^{T}\otimes\mathbf{I}_{\bar{M}})\mathbf{B}\mathbf{P}$, and $\mathbf{d}_{g}(\bar{\bm{\phi}}_{g})\triangleq\mathsf{vec}(Y_{0}\mathbf{I}_{\bar{M}}-\overline{\mathsf{vec}}(\mathbf{P}\bar{\bm{\phi}}_{g}))$. In the following ADMM framework, (23) will be employed for updating the variable $\bar{\bm{\phi}}$, while (24) will be used for updating the variable $\bar{\mathbf{y}}_{g}$. However, updating $\bar{\mathbf{y}}_{g}$ meets a challenge due to the non-separable nature between entries of $\bar{\mathbf{y}}_{g}$ in (24), which makes it impossible to individually optimize each entry in $\bar{\mathbf{y}}_{g}$ under the non-convex constraint (17d). To overcome this, one effective solution is to introduce an auxiliary variable $\mathbf{z}$ by setting $\mathbf{z}=\bar{\mathbf{y}}$, where $\bar{\mathbf{y}}=[\bar{\mathbf{y}}_{1}^{T},\ldots,\bar{\mathbf{y}}_{G}^{T}]^{T}$, thereby decoupling (24) from the troublesome constraint (17d). Hence, problem (19) is equivalently re-written as
@@
$[\mathbf{z}]_{\tilde{m}}\in\mathcal{Y},~\forall\tilde{m}\in\tilde{\mathcal{M}}$, (25d)
where $\tilde{\mathcal{M}}=\{1,\ldots,GU\}$. Following the ADMM framework, we construct the augmented Lagrangian (AL) function in its scaled form [boyd_admm], given by
@@
where $\mathbf{C}=\mathsf{blkdiag}(\mathbf{C}_{1},\ldots,\mathbf{C}_{G})$, $\mathbf{c}=[\mathbf{c}_{1}^{T},\ldots,\mathbf{c}_{G}^{T}]^{T}$, and $\bm{\lambda}_{1}\in\mathbb{C}^{G\bar{M}^{2}\times 1}$, $\bm{\lambda}_{2}\in\mathbb{C}^{U\times 1}$ denote the scaled Lagrange multipliers corresponding to constraints (25b) and (25c), respectively, and $\rho_{1}$ and $\rho_{2}$ are penalty coefficients.
@@
    $\bar{\bm{\phi}}=\left(\rho_{1}\mathbf{C}^{H}\mathbf{C}\right)^{-1}\left(\bar{\mathbf{f}}_{t}+\rho_{1}\mathbf{C}^{H}(\mathbf{c}-\bm{\lambda}_{1})\right);$ (28)
@@
    $\bar{\mathbf{y}}=\Big(\rho_{1}\mathbf{D}^{H}\mathbf{D}$ $+\rho_{2}\mathbf{I}_{GU}\Big)^{-1}$ (30)
    $\times\left(\rho_{1}\mathbf{D}^{H}(\mathbf{d}-\bm{\lambda}_{1})+\rho_{2}(\mathbf{z}+\bm{\lambda}_{2})\right);$
@@
    Since constraint (25d) is separable with respect to each entry of vector $\mathbf{z}$, problem (31) can be decomposed into GUGU independent one-dimensional problems. Defining $\mathbf{r}\triangleq\bar{\mathbf{y}}-\bm{\lambda}_{2}$, each one-dimensional problem can be expressed as
@@
[…]
@@
Algorithm 1 Proposed $M$-ADMM Algorithm
[…]
@@
4:while no convergence of objective (17a) do
5:  while no convergence of objective (19a) do
@@
8:   Update $\mathbf{z}]_{\tilde{m}},~\forall\tilde{m}\in\tilde{\mathcal{M}}$ by (34).
@@ III-B.3 Summary
The proposed $M$-ADMM algorithm is summarized in Algorithm 1. Once $\bar{\bm{\phi}}^{\star}$ is obtained, we can compute $\bm{\Phi}^{\star}_{g}$ as $\bm{\Phi}^{\star}_{g}=\overline{\mathsf{vec}}(\mathbf{P}\bar{\bm{\phi}}^{\star}_{g}),~\forall g\in\mathcal{G}$. The final solution for $\bm{\Phi}^{\star}$ is then given by $\bm{\Phi}^{\star}=\mathsf{blkdiag}(\bm{\Phi}^{\star}_{1},\ldots,\bm{\Phi}^{\star}_{G})$.
@@ III-C Low-Complexity Algorithm
The proposed $M$-ADMM algorithm requires a double-loop iteration, which might require high computational complexity to guarantee convergence. To ease the requirement of the double-loop iteration, we first recall that, for a lossless group-connected reconfigurable admittance network, the received signal power $\left|h_{RT}+\mathbf{h}_{RI}^{H}\bm{\Phi}\mathbf{h}_{IT}\right|^{2}$ reaches its upper bound when $\bm{\Phi}$ satisfies [shen2021modelling]
$e^{\jmath\arg(h_{RT})}\bar{\mathbf{h}}_{RI,g}=\bm{\Phi}_{g}\bar{\mathbf{h}}_{IT,g},~\forall g\in\mathcal{G},$ (37)
where $\bar{\mathbf{h}}_{RI}=[\bar{\mathbf{h}}_{RI,1}^{T},\ldots,\bar{\mathbf{h}}_{RI,G}^{T}]^{T}$ with $\bar{\mathbf{h}}_{RI,g}\triangleq\frac{\mathbf{h}_{RI,g}}{\|\mathbf{h}_{RI,g}\|_{2}}\in\mathbb{C}^{\bar{M}\times 1}$, and $\bar{\mathbf{h}}_{IT}=[\bar{\mathbf{h}}_{IT,1}^{T},\ldots,\bar{\mathbf{h}}_{IT,G}^{T}]^{T}$ with $\bar{\mathbf{h}}_{IT,g}\triangleq\frac{\mathbf{h}_{IT,g}}{\|\mathbf{h}_{IT,g}\|_{2}}\in\mathbb{C}^{\bar{M}\times 1}$. By substituting $\bm{\Phi}_{g}$ with $\mathbf{Y}_{g},~\forall g\in\mathcal{G}$ based on (17b), (37) can be equivalently re-written as
@@
$\mathrm{s.t.}$~~ $[\bar{\mathbf{y}}]_{\tilde{m}}\in\mathcal{Y},~\forall\tilde{m}\in\tilde{\mathcal{M}}$, (40b)
where $\mathbf{A}\in\mathbb{C}^{M\times GU}$ is a block-diagonal matrix given by $\mathbf{A}=\mathsf{blkdiag}(\mathbf{A}_{1},\ldots,\mathbf{A}_{G})$, with $\mathbf{A}_{g}\in\mathbb{C}^{\bar{M}\times U},~\forall g\in\mathcal{G}$ derived from vector $\mathbf{a}_{g}$ through a series of transformations. Each $\mathbf{a}_{g}\in\mathbb{C}^{\bar{M}\times 1}$ represents a block of the vector $\mathbf{a}$, which is defined as $\mathbf{a}=[\bar{\mathbf{a}}^{T}_{1},\ldots,\bar{\mathbf{a}}^{T}_{G}]^{T}$. To prove the transformation in (40a), we utilize the vectorization property to reformulate $\mathbf{Y}_{g}\mathbf{a}_{g},~\forall g\in\mathcal{G}$ as
@@
$\mathbf{Y}_{g}\mathbf{a}_{g}=\mathbf{A}_{g}\bar{\mathbf{y}}_{g},~\forall g\in\mathcal{G},$ (42)
@@
We apply the ADMM method [boyd_admm] to solve (43). The AL function of (43) is given by:
@@
where $\mathbf{u}\in\mathbb{C}^{GU\times 1}$ is the scaled Lagrangian multiplier and $\rho$ is the penalty coefficient. The detailed algorithm is presented below.
@@
    $\bar{\mathbf{y}}=\left(2\mathbf{A}^{H}\mathbf{A}+\rho\mathbf{I}_{GU}\right)^{-1}\left(2\mathbf{A}^{H}\mathbf{b}+\rho(\mathbf{z}+\mathbf{u})\right);$ (46)
@@
With the solution $\bar{\mathbf{y}}^{\star}$ to problem (43), we can compute $\mathbf{Y}_{g}^{\star},~\forall g\in\mathcal{G}$ as $\mathbf{Y}_{g}^{\star}=\overline{\mathsf{vec}}(\mathbf{B}\mathbf{P}\bar{\mathbf{y}}_{g}^{\star})$. The admittance matrix is then given by $\mathbf{Y}^{\star}=\mathsf{blkdiag}(\mathbf{Y}_{1}^{\star},\ldots,\mathbf{Y}_{G}^{\star})$, from which the scattering matrix $\bm{\Phi}^{\star}$ is obtained using (1).
Remark 3: The only difference between group-connected and forest-connected architectures lies in the number of variables in each $\bar{\mathbf{Y}}_{g}$. For the group-connected BD-RIS, $\bar{\mathbf{Y}}_{g},~\forall g\in\mathcal{G}$ is a matrix with no zero entries, whereas certain entries in $\bar{\mathbf{Y}}_{g}$ are zero according to (5) or (6) for forest-connected BD-RIS. This difference affects our optimization algorithms only through the mapping $\bar{\mathbf{Y}}_{g}=\overline{\mathsf{vec}}(\mathbf{P}\bar{\mathbf{y}}_{g})$, where matrix $\mathbf{P}$ differs for the group-connected and forest-connected BD-RISs, as specified in (62) – (64); each $\bar{\mathbf{y}}_{g}$ contains $\frac{\bar{M}(\bar{M}+1)}{2}$ entries for the group-connected BD-RIS and $2\bar{M}-1$ entries for the forest-connected BD-RIS. Apart from this modification, both the $M$-ADMM and the low-complexity algorithm apply in the same way to the two architectures.
III-D Complexity Analysis